OpenStudy (anonymous):

how do i solve the systems 2m + 4n = 10 3m + 5n = 11 and 5x + y = 0 5x + 2y = 30 using elimination?

OpenStudy (anonymous):

2m + 4n = 10 -- (I) 3m + 5n = 11 -- (II) Multiply (I) by 3 and (II) by -2 nd then add the resulting equations. This way m will be eliminated and the remaining equation will give value of n. Then put this value of n in either (I) or (II) and get the value of m

OpenStudy (anonymous):

5x + y = 0 --(I) 5x + 2y = 30 --(II) Multiply (I) by 1 and (II) by -1 and add. x will be eliminated and from the resulting equation you can find the value of y. Then put that value in either (I) or (II) and get the value of x